Wendy, a loan officer at a bank, has $1,000,000 to lend and is required to obtain an average return of 8% per year. If she can lend at the rate of 7% or at the rate of 11%, how much can she lend at the 7% rate and still meet her requirement?
Answer by nyc_function(2741)

We use the simple interest formula I = Prt where I = interest earned, P = principal invested, r = rate, and t = time
We want an I of (1000000 * 0.08) = $80,000
let x = amount loaned at 7%
(1000000 - x) = amount loaned at 11%
So...
80000 = x(0.07)(1) + (1000000 - x)(0.11)(1)
80000 = 0.07x + 110000 - 0.11x
.04x = 30000
x = 30000/0.04
x = $750,000
To check:
750000 * 0.07 = 52500
250000 * 0.11 = 27500
52500 + 27500 = 80000
I hope this helps.